Find the value of T if:
P+Q=R
R+S=T
P+T=U
Q+S+U= 40

Answer by solver91311(24713)



Q + S + U = 40

But P + T = U, so Q + S + P + T = 40

But P + Q = R, so S + R + T = 40

But S + R = T, so T + T = 40.

You can finish this, I think.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism